A vertically placed ideal spring at the lower end of which the body has an oscillation period of 7.2 s. How long will the mentioned system have on the Moon, if we know that the gravitational acceleration on the Moon is 6 times less than that on the Earth's surface? (Answer: T = 7.2 s)


1
Expert's answer
2021-04-13T06:41:11-0400

According to the equation for period of a mass-spring system:


"T=2\pi\sqrt{\frac{m}{k}}."

This equation does not contain acceleration due to gravity, the period depends on the mass and spring rate only. So, it will not change.